OpenStudy (anonymous):

The order of a chemical reaction with respect to reactant A is zero. If you double [A], the rate a) remains the same b) doubles c) decreases by half d) quadruples

OpenStudy (unklerhaukus):

\[-\frac{\text d[A]}{\text dt}=k[A]^0\] \[-\frac{\text d[A]}{\text dt}=k\]

OpenStudy (anonymous):

So the answer would be B. A would be doubled.

OpenStudy (unklerhaukus):

how did you get that

OpenStudy (anonymous):

Wait no

OpenStudy (anonymous):

It would stay the same

OpenStudy (unklerhaukus):

\[-\frac{\text d[A]}{\text dt}=k\] double \([A]\) \[-\frac{\text d[A]}{\text dt}=k\] the rate is still \(k\)

OpenStudy (anonymous):

since A is zero it would stay the same. [A]^0

OpenStudy (unklerhaukus):

i think you mean \[[A]^0=[2A]^0=1\]

OpenStudy (anonymous):

Yeah, accidentally submitted it without finishing it

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

Appreciate your patience

OpenStudy (unklerhaukus):

any time